# Energetically consistent localised APE budgets for local and regional studies of stratified flow energetics

### v1

@@ 2.2 Local APE theory
$E_{a}(\rho,z)=\int_{\rho^{{\ddagger}}}^{\rho}\frac{\partial E_{a}}{\partial\rho}(\tilde{\rho},z)\,{\rm d}\tilde{\rho}=\frac{g}{\rho_{\star}}\int_{\rho^{{\ddagger}}}^{\rho}[z-z_{0}(\tilde{\rho})]\,{\rm d}\tilde{\rho},$ (13)
@@ 2.4 Convexity of APE density and eddy APE
Physically, the non-negative character of $N_{0}^{2}$ in Eq. 16) follows from Lorenz reference state being a state of minimum potential energy, hence statically stable by construction, which establishes convexity with respect to z. To prove the non-negative character of (15) and hence the convexity with respect to density, simply differentiate the the LNB equation (12) with respect to $\rho$, thus leading to
@@
$\begin{split}A_{e}=&E_{a}(\rho,z)-E_{a}(\overline{\rho},z)-\frac{\partial E_{a}}{\partial\rho}(\overline{\rho},z)(\rho-\overline{\rho})\\ =&\int_{\overline{\rho}}^{\rho}\left[\frac{\partial E_{a}}{\partial\rho}(\tilde{\rho},z)-\frac{\partial E_{a}}{\partial\rho}(\overline{\rho},z)\right]\,{\rm d}\tilde{\rho}=-\frac{g}{\rho_{\star}}\int_{\overline{\rho}}^{\rho}[z_{0}(\tilde{\rho})-z_{0}(\overline{\rho})]\,{\rm d}\tilde{\rho}\end{split}$ (36)
@@
The approximation (38) is identical to the expression used by Roullet et al. (2014). However, while the buoyancy anomaly $b^{\prime}$ is the same as in Roullet’s approach, the displacement $\zeta^{\prime}$ is defined quite differently in terms of the instantaneous and mean reference positions $z_{0}(\rho)$ and $z_{0}(\overline{\rho})$ respectively, that is, in terms of the equilibrium positions of $\rho$ and $\overline{\rho}$ in Lorenz reference state.
@@
where $\Lambda$ is the same parameter introduced previously. This establishes that the validity and accuracy of the heuristic localised APE density depend on the proximity of the actual state to Lorenz reference state, which does not appear to have been previously recognised.
@@ 3.3 Eddy APE budget and turbulent APE dissipation
$\kappa\nabla\overline{\rho}\cdot\nabla(\overline{\Upsilon}-\Upsilon_{m})\approx\frac{g}{\rho_{\star}}\frac{\partial^{2}z_{0}}{\partial\rho}(\overline{\rho})\kappa\nabla\overline{\rho}\cdot\nabla\frac{\overline{\rho^{\prime 2}}}{2}+\frac{g}{\rho_{\star}}\frac{\overline{\rho^{\prime 2}}}{2}\frac{\partial^{3}z_{0}}{\partial\rho^{3}}(\overline{\rho})\kappa|\nabla\overline{\rho}|^{2}$ (67)
@@
$\varepsilon_{p}^{t}\approx-\frac{g}{\rho_{\star}}\frac{\partial z_{0}}{\partial\rho}(\overline{\rho})\kappa\overline{|\nabla\rho^{\prime}|^{2}}-\frac{g}{\rho_{\star}}\frac{\overline{\rho^{\prime 2}}}{2}\frac{\partial^{3}z_{0}}{\partial\rho^{3}}(\overline{\rho})\kappa|\nabla\overline{\rho}|^{2}$ (69)
@@ 4 Application to energetically consistent numerical ocean modelling
Our framework provides rigorous theoretical foundations for linking parameterised energy transfers to observable KE and APE dissipation rates. In numerical ocean models, the main energy transfers of interest are those associated with the ocean meso-scale, responsible for the transfer of mean APE to turbulent or eddy kinetic energy, which can then be subsequently dissipated through irreversible viscous processes, and those associated with small-scale turbulent stratified mixing, responsible for the transfer of turbulent kinetic energy into turbulent APE, which may then be dissipated by irreversible diffusive processe giving rise to turbulent diapycnal mixing. The following aims to illustrate these ideas.
@@ 4.1 Energetically consistent modelling
Parameterisations of subgridscale processes control the energy transfers between the resolved and unresolved scales of motions implicated in the turbulent forward KE and APE energy cascades. Upon reaching molecular scales, mechanical energy can be dissipated quasi-adiabatically as KE at the viscous dissipation rate $\varepsilon_{k}^{t}$, or diabatically as APE at the diffusive dissipation rate $\varepsilon_{p}^{t}$. A key difficulty is ensuring the consistency of the parameterised energy transfers with the constraints on the dissipation ratio $\Gamma=\varepsilon_{p}^{t}/\varepsilon_{k}^{t}$, which is traditionally regarded as a measure of mixing efficiency close to $0.2$. Addressing this issue is the main objective of energetically consistent modelling approaches proposed by Carsten Eden and his group (Eden et al., 2014; Eden, 2015, 2016). To that end, the prevailing approach has relied on using a turbulent kientic energy (TKE) equation to predict the turbulent viscous dissipation $\varepsilon_{k}^{t}$, from which the APE dissipation rate $\varepsilon_{p}^{t}$ and turbulent vertical mixing coefficient $K_{\rho}=\varepsilon_{p}^{t}/N^{2}$ may be inferred, provided that the value of the dissipation ratio $\Gamma$ can also be predicted in some way. For a review of our current understanding of $\Gamma$, see Gregg (2021). An alternative that has yet to be developed in oceanography would be to predict $\varepsilon_{p}^{t}$ directly from a turbulent APE equation, as proposed in the context of atmospheric boundary layer research by Zilitinkevich and collaborators, e.g., Zilitinkevich et al. (2013). An important advantage of the eddy APE budget is that it constraints the full turbulent density flux $\overline{\rho^{\prime}{\bf v}^{\prime}}$, whereas the TKE budget depends only on the vertical component $\overline{\rho^{\prime}w^{\prime}}$.
@@
To analyse inter-scale energy transfers, a filtering approach is needed to isolate the different scales of interest. In the context of large-scale ocean modelling, Eden (2015) proposed that the subgrid-scale energy should be divided into subreservoirs for meso-scale eddies, internal gravity waves, and turbulent incoherent motions. Here, we only consider a subdivision of energy into mean (resolved) and eddy (unresolved) scales due to the inherent limitations of standard Reynolds averaging. With this in mind, we return to the eddy APE budget (46), which we rewrite in the form
@@
where $C(E_{k}^{t},E_{a}^{t})$ and $C(E_{a}^{m},E_{a}^{t})$ are the conversions of turbulent kinetic energy and mean APE into eddy APE, respectively, whose expressions are
@@
Under the classical assumptions of stationarity and homogeneity, the terms in the left-hand side of (70) can be neglected, and the eddy APE budget reduces to a balance between production of eddy APE due to conversions with mean APE and eddy KE and turbulent APE dissipation, viz.,
@@
which from (71) and (72) may be written as
@@
Eq. (74) can be understood as a constraint on the diapycnal component of the turbulent density flux across the mean isopycnal surfaces $\overline{\rho}={\rm constant}$, controlled by the turbulent APE dissipation rate $\varepsilon_{p}^{t}$. In the case where the resolved flow pertains to the large scales, the turbulent density flux is generally assumed to contain at least two components
@@
pertaining to the effects of meso-scale eddies and small scale turbulent mixing, respectively, being taken to be perpendicular and parallel to $\nabla\overline{\rho}$ according to
@@
e.g., Griffies (1998); Griffies et al. (1998). Of particular interest is the vertical component
@@ 4.2 Energetics of downgradient diffusion
e.g., Lindborg and Brethouwer (2008), where $\Gamma$ is the dissipation ratio Oakey (1982). However, according to the eddy APE budget (76), a more accurate expression is
@@
which may be re-arranged under the form
@@
in which $\Lambda$ and ${\bf S}$ are defined as before by (26). As stated earlier, Eq. (80) shows that the standard expression (78) implicitly depends on two assumptions that are rarely, if ever, acknowledged: 1) that it depends on the local stratification as measured by $\overline{N}^{2}$ being approximately equal to $N_{0}^{2}$; 2) the small slope approximation $|{\bf S}|\ll 1$ that is often made in the context of the rotated Redi diffusion (Redi, 1982) for instance. Moreover, if we neglect the higher order terms in the definition of the eddy APE dissipation rate so that
@@
Physically, Eq. (82) states that the resolved dissipation of the mean density field must be ultimately balanced by the dissipation of eddy density variance at molecular scales. Note, however, that when approached from the eddy APE budget viewpoint, (82) requires neglecting the higher order terms in the eddy APE dissipation rate. In the literature, (82) is more commonly obtained by equating the production of density variance by the turbulent density flux with the dissipation of density variance, as in the Osborn-Cox model (Osborn and Cox, 1972). Note that Eq. (82) cannot give rise to upgradient (negative) diffusion, while (80) can potentially allow it in some circumstances.
@@
In the literature, the mean to eddy APE conversion has been exclusively discussed in the context of the quasi-geostrophic (QG) approximation, e.g., von Storch et al. (2012), which is equivalent to assuming $\Lambda\approx 1$ in (83), in which case it reduces to
@@
and always acts as a downscale transfer of energy from mean APE to eddy APE. The exact finite-amplitude mean APE to eddy APE conversion may occasionally behave quite differently, as Eq. (83) shows that the sign of the conversion is no longer necessarily non-negative. Energy transfer can be either upscale or downscale, depending on the relative values of $\Lambda$ and of the slope parameter $|{\bf S}|$. Physically, the condition for upscale transfer from eddy APE to mean APE to occur is that $\Lambda<1$ such that
$\Lambda<\frac{1}{1+|{\bf S}|^{2}}\qquad{\rm or}\qquad|{\bf S}|^{2}<\frac{1}{\Lambda}-1\quad{\rm if}\quad\Lambda<1$ (86)
and therefore may also be regarded as a criterion on the slope vector. In that case, downgradient diffusion can backscatter unresolved energy into resolved energy, at least in principle, thus potentially acting as a source of instability for the resolved flow. Although the associated upscale energy transfer will be counteracted by the downscale energy transfer associated with skew diffusion, it is not necessarily obvious that this can actually suppress the diffusive instability, because downgradient diffusion and skew diffusion a priori operate differently.
@@ 4.3 Energetics of skew diffusion and TKE budget
Physically, skew-diffusion can be interpreted as an eddy-induced advection by subgrid-scale processes, as follows from the relation
@@
As a result, the evolution equation for the mean density $\overline{\rho}$ may be written in the form
@@
with ${\bf v}_{res}=\overline{\bf v}+{\bf v}_{eddy}$ being the residual velocity and ${\bf v}_{eddy}=\nabla\times\boldsymbol{\Psi}$ the eddy-induced velocity. By construction, the eddy induced velocity is divergenceless $\nabla\cdot{\bf v}_{eddy}=0$. In oceanography, the most commonly used parameterisation for $\boldsymbol{\Psi}$ is
@@
Of particular interest is the vertical component, which is given by
@@
which acts as ‘upgradient’ diffusion.
@@
$\left\{C(E_{a}^{m},E_{a}^{t})\right\}_{skew}=-\overline{\rho^{\prime}{\bf v}^{\prime}}_{skew}\cdot\nabla\Upsilon_{m}=-\frac{g}{\rho_{\star}}\overline{\rho^{\prime}w^{\prime}}_{skew}=\kappa_{gm}|{\bf S}|^{2}\overline{N}^{2}$ (92) $\left\{C(E_{k}^{t},E_{a}^{t})\right\}_{skew}=\frac{g}{\rho_{\star}}\overline{\rho^{\prime}w^{\prime}}_{skew}=-\kappa_{gm}|{\bf S}|^{2}\overline{N}^{2}$ (93)
Eq. (92) confirms that skew diffusion acts as a net sink of mean APE whose magnitude is proportional to the slope squared $|{\bf S}|^{2}$ and mean squared buoyancy frequency $\overline{N}^{2}$.

### v2

@@ 2.2 Local APE theory
$E_{a}(\rho,z)=\int_{\rho^{{\ddagger}}(z)}^{\rho}\frac{\partial E_{a}}{\partial\rho}(\tilde{\rho},z)\,{\rm d}\tilde{\rho}=\frac{g}{\rho_{\star}}\int_{\rho^{{\ddagger}}(z)}^{\rho}[z-z_{0}(\tilde{\rho})]\,{\rm d}\tilde{\rho},$ (13)
@@ 2.4 Convexity of APE density and eddy APE
Physically, the non-negative character of $N_{0}^{2}$ in Eq. 16) follows from Lorenz reference state being a state of minimum potential energy, hence statically stable by construction, which establishes convexity with respect to $z$. To prove the non-negative character of (15) and hence the convexity with respect to density, simply differentiate the LNB equation (12) with respect to $\rho$, thus leading to
@@
$\begin{split}A_{e}=&E_{a}(\rho,z)-E_{a}(\overline{\rho},z)-\frac{\partial E_{a}}{\partial\rho}(\overline{\rho},z)(\rho-\overline{\rho})\\ =&\int_{\overline{\rho}}^{\rho}\left[\frac{\partial E_{a}}{\partial\rho}(\tilde{\rho},z)-\frac{\partial E_{a}}{\partial\rho}(\overline{\rho},z)\right]\,{\rm d}\tilde{\rho}\\ =&-\frac{g}{\rho_{\star}}\int_{\overline{\rho}}^{\rho}[z_{0}(\tilde{\rho})-z_{0}(\overline{\rho})]\,{\rm d}\tilde{\rho}\end{split}$ (36)
@@
The approximation (38) is identical to the expression used by Roullet et al. (2014). However, while the buoyancy anomaly $b^{\prime}$ is the same as in Roullet’s approach, the displacement $\zeta^{\prime}$ is defined quite differently in terms of the instantaneous and mean reference positions $z_{0}(\rho)$ and $z_{0}(\overline{\rho})$ respectively, that is, in terms of the equilibrium positions of $\rho$ and $\overline{\rho}$ in Lorenz reference state. These differences are clearly illustrated in the right panel of Fig. 1 introduced earlier.
@@
where $\Lambda$ is the same parameter introduced previously. This establishes that the validity and accuracy of the heuristic localised APE density depend on the proximity of the actual state to Lorenz reference state, as expected.
@@ 3.3 Eddy APE budget and turbulent APE dissipation
$\kappa\nabla\overline{\rho}\cdot\nabla(\overline{\Upsilon}-\Upsilon_{m})\approx\frac{g}{\rho_{\star}}\frac{\partial^{2}z_{0}}{\partial\rho^{2}}(\overline{\rho})\kappa\nabla\overline{\rho}\cdot\nabla\frac{\overline{\rho^{\prime 2}}}{2}+\frac{g}{\rho_{\star}}\frac{\overline{\rho^{\prime 2}}}{2}\frac{\partial^{3}z_{0}}{\partial\rho^{3}}(\overline{\rho})\kappa|\nabla\overline{\rho}|^{2}$ (67)
@@
$\varepsilon_{p}^{t}\approx-\frac{g}{\rho_{\star}}\frac{\partial z_{0}}{\partial\rho}(\overline{\rho})\kappa\overline{|\nabla\rho^{\prime}|^{2}}{\color[rgb]{0,0,0}+}\frac{g}{\rho_{\star}}\frac{\overline{\rho^{\prime 2}}}{2}\frac{\partial^{3}z_{0}}{\partial\rho^{3}}(\overline{\rho})\kappa|\nabla\overline{\rho}|^{2}$ (69)
@@ 4 Application to energetically consistent numerical ocean modelling
Our framework provides a rigorous theoretical foundation for linking parameterised energy transfers to observable KE and APE dissipation rates. In numerical ocean models, the main energy transfers of interest are those associated with the ocean meso-scale, which are responsible for transferring mean APE to turbulent or eddy kinetic energy (which is subsequently dissipated through irreversible viscous processes), and those associated with small-scale turbulent stratified mixing, which transfer turbulent kinetic energy into turbulent APE, later dissipated by irreversible diffusive processes, giving rise to turbulent diapycnal mixing. The following aims to illustrate these ideas.
@@ 4.1 Energetically consistent modelling
Parameterisations of subgridscale processes control the energy transfers between the resolved and unresolved scales of motion implicated in the turbulent forward KE and APE energy cascades. Upon reaching molecular scales, mechanical energy can be dissipated quasi-adiabatically as KE at the viscous dissipation rate $\varepsilon_{k}^{t}$, or diabatically as APE at the diffusive dissipation rate $\varepsilon_{p}^{t}$. In most numerical ocean models, subgrid-scale process parameterisations have traditionally been developed without explicit consideration of their implications for the KE and APE cascades or for the KE and APE dissipation rates. However, this changed when Munk and Wunsch (1998) noted that the intensity of mixing depends on the intensity of the mechanical sources of stirring. In particular, they raised the issue of whether tuning the value of the turbulent diapycnal or vertical mixing coefficients $K_{\rho}$ or $K_{v}$ to minimise mismatch with observed temperature and salinity fields was necessarily consistent with our understanding of how turbulent mixing is mechanically sustained. To that end, they considered the gravitational potential energy (GPE) budget, assuming a balance between the rate of GPE increase due to mixing at the rate $K_{v}N^{2}$ and the rate of GPE loss due to high-latitude cooling (neglecting the role of nonlinearities in the equation of state). To link $K_{v}$ to the mechanical sources of stirring, they used the relation $K_{v}N^{2}=\Gamma\varepsilon_{K}$ with $\Gamma=0.2$, and balanced the total viscous dissipation rate with the power input due to winds and tides, dismissing the role of surface buoyancy fluxes (see Tailleux (2010) for a discussion). The work of Munk and Wunsch (1998) proved very influential, prompting the development of ‘energetically consistent modelling’ pioneered by Carsten Eden and his group (Eden et al., 2014; Eden, 2015, 2016), whose aim is to link the intensity of mixing to the sources of stirring by exploiting empirical knowledge about the dissipation ratio $\Gamma=\varepsilon_{p}/\varepsilon_{k}$. In our view, this aim can be more generally defined as seeking to exploit existing and developing knowledge about the inter-relations between the KE and APE budgets, for instance by developing and using a turbulent APE budget in addition to the widely used turbulent KE budget.
@@
In practice, the prevailing approach has primarily relied on using a turbulent kinetic energy (TKE) equation to predict the turbulent viscous dissipation $\varepsilon_{k}^{t}$, from which the APE dissipation rate $\varepsilon_{p}^{t}$ and turbulent vertical mixing coefficient $K_{\rho}=\varepsilon_{p}^{t}/N^{2}$ can be inferred, provided that the value of the dissipation ratio $\Gamma$ can also be predicted in some way. For a review of our current understanding of $\Gamma$, see Gregg (2021). An alternative, yet to be developed in oceanography, would be to predict $\varepsilon_{p}^{t}$ directly from a turbulent APE equation, as proposed in the context of atmospheric boundary layer research by Zilitinkevich and collaborators (e.g., Zilitinkevich et al. (2013)). An important advantage of the eddy APE budget is that it constrains the full turbulent density flux $\overline{\rho^{\prime}{\bf v}^{\prime}}$, whereas the TKE budget depends only on the vertical component $\overline{\rho^{\prime}w^{\prime}}$, which is why we think it should receive more attention.
@@
To analyse across-scale energy transfers, a filtering approach is needed to isolate the different scales of interest. In the context of large-scale ocean modelling, Eden (2015) proposed that the subgrid-scale energy should be divided into subreservoirs for meso-scale eddies, internal gravity waves, and turbulent incoherent motions. Here, we only consider a subdivision of energy into mean (resolved) and eddy (unresolved) scales due to the inherent limitations of standard Reynolds averaging. With this in mind, we return to the eddy APE budget (46), which we rewrite as
@@
where $C(E_{k}^{t},E_{a}^{t})$ and $C(E_{a}^{m},E_{a}^{t})$ are the conversions of turbulent kinetic energy and mean APE into eddy APE, respectively, with
@@
Under the classical assumptions of stationarity and homogeneity, the terms on the left-hand side of (70) can be neglected, and the eddy APE budget reduces to a balance between production of eddy APE due to conversions with mean APE and eddy KE, and turbulent APE dissipation, viz.,
@@
which, from (71) and (72), may be written as
@@
Eq. (74) can be understood as a constraint on the diapycnal component of the turbulent density flux across the mean isopycnal surfaces $\overline{\rho}={\rm constant}$, controlled by the turbulent APE dissipation rate $\varepsilon_{p}^{t}$. When the resolved flow pertains to large scales, the turbulent density flux is generally assumed to contain at least two components:
@@
pertaining to the effects of meso-scale eddies and small-scale turbulent mixing, respectively, and are taken to be perpendicular and parallel to $\nabla\overline{\rho}$ according to
@@
(e.g., Griffies (1998); Griffies et al. (1998)). Of particular interest is the vertical component
@@ 4.2 Energetics of downgradient diffusion
(e.g., Lindborg and Brethouwer (2008)), where $\Gamma$ is the dissipation ratio Oakey (1982). However, according to the eddy APE budget (76), a more accurate expression is
@@
which may be rearranged as
@@
where $\Lambda$ and ${\bf S}$ are defined as before by (26). As stated earlier, Eq. (80) shows that the standard expression (78) implicitly depends on two assumptions that are rarely, if ever, acknowledged: 1) that the local stratification as measured by $\overline{N}^{2}$ is approximately equal to $N_{0}^{2}$; and 2) the small slope approximation $|{\bf S}|\ll 1$, often made in the context of rotated Redi diffusion (Redi, 1982), for instance. Moreover, if we neglect the higher order terms in the definition of the eddy APE dissipation rate so that
@@
Physically, Eq. (82) states that the resolved dissipation of the mean density field must ultimately be balanced by the dissipation of eddy density variance at molecular scales. Note, however, that when approached from the eddy APE budget viewpoint, (82) requires neglecting the higher order terms in the eddy APE dissipation rate. In the literature, (82) is more commonly obtained by equating the production of density variance by the turbulent density flux with the dissipation of density variance, as in the Osborn-Cox model (Osborn and Cox, 1972). Note that Eq. (82) cannot give rise to upgradient (negative) diffusion, while (80) can potentially allow it in some circumstances.
@@
In the literature, the mean to eddy APE conversion has been exclusively discussed in the context of the quasi-geostrophic (QG) approximation (e.g., von Storch et al. (2012)), which is equivalent to assuming $\Lambda\approx 1$ in (83), in which case it reduces to
@@
and always acts as a downscale transfer of energy from mean APE to eddy APE. The exact finite-amplitude mean APE to eddy APE conversion may occasionally behave quite differently, as Eq. (83) shows that the sign of the conversion is no longer necessarily non-negative. Energy transfer can be either upscale or downscale, depending on the relative values of $\Lambda$ and the slope parameter $|{\bf S}|$. Because the value of $\Lambda$ is controlled by the distance from the Lorenz reference state, which itself depends on the steepness of isopycnal slopes, we assume that the condition for upscale transfer from eddy APE to mean APE imposes a constraint on $\Lambda$ (rather than on $|{\bf S}|$), namely
$\Lambda<\frac{1}{1+|{\bf S}|^{2}}.$ (86)
In that case, downgradient diffusion can, at least in principle, backscatter unresolved energy into resolved energy, thus potentially acting as a source of instability for the resolved flow. Although the associated upscale energy transfer will be counteracted by the downscale energy transfer associated with skew diffusion, it is not necessarily obvious that this can actually suppress the diffusive instability, because downgradient diffusion and skew diffusion a priori operate differently.
@@ 4.3 Energetics of skew diffusion and TKE budget
Physically, skew diffusion can be interpreted as an eddy-induced advection by subgrid-scale processes, as follows from the relation
@@
As a result, the evolution equation for the mean density $\overline{\rho}$ may be written as
@@
with ${\bf v}_{res}=\overline{\bf v}+{\bf v}_{eddy}$ being the residual velocity and ${\bf v}_{eddy}=\nabla\times\boldsymbol{\Psi}$ the eddy-induced velocity. By construction, the eddy-induced velocity is divergenceless, $\nabla\cdot{\bf v}_{eddy}=0$. In oceanography, the most commonly used parameterisation for $\boldsymbol{\Psi}$ is
@@
Of particular interest is the vertical component, given by
@@
which acts as an ‘upgradient’ flux.
@@
$\begin{split}\left\{C(E_{a}^{m},E_{a}^{t})\right\}_{skew}=&-\overline{\rho^{\prime}{\bf v}^{\prime}}_{skew}\cdot\nabla\Upsilon_{m}\\ =&-\frac{g}{\rho_{\star}}\overline{\rho^{\prime}w^{\prime}}_{skew}=\kappa_{gm}|{\bf S}|^{2}\overline{N}^{2}\end{split}$ (92) $\left\{C(E_{k}^{t},E_{a}^{t})\right\}_{skew}=\frac{g}{\rho_{\star}}\overline{\rho^{\prime}w^{\prime}}_{skew}=-\kappa_{gm}|{\bf S}|^{2}\overline{N}^{2}$ (93)
Eq. (92) confirms that skew diffusion acts as a net sink of mean APE whose magnitude is proportional to the slope squared $|{\bf S}|^{2}$ and mean squared buoyancy frequency $\overline{N}^{2}$ e.g., Griffies (2004).